HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

LINE-BY-LINE INTENSITY MEASUREMENTS OF METHYL FLUORIDE IN THE ν_{2} and ν_{5} BANDS

\underline{H. ZIADI}, \underline{O. BEN FATHALLAH} , \underline{I. BEN CHOUIKHA} , Laboratory of

Spectroscopy and Molecular Dynamics, National Superior School of Engineering of

Tunis, 5 Av Taha Hussein, 1008 Tunis, Tunisia; \underline{F. KWABIA TCHANA} , \underline{X. LANDSHEERE} , Université Paris Cité and Univ Paris Est Creteil, CNRS, LISA,

F-75013 Paris, France ; and \underline{H. AROUI}, Laboratory of Spectroscopy and Molecular Dynamics, National Superior School of Engineering of Tunis, 5 Av Taha

Hussein, 1008 Tunis, Tunisia

Systematic measurements of line intensities in the ν_{2} and ν_{5} bands of CH_{3}F were performed for the first time for 2040 transitions using a mono-spectrum fitting technique utilizing Voigt and Rautian profiles. The spectra are recorded at room temperature for eight pressures of CH_{3}F from 0.1247 to 2.5080 Torr using the Bruker IFS125HR Fourier Transform spectrometer located at the LISA facility in Créteil. Strong intensity perturbations were observed and analyzed, accounting for the strong Coriolis coupling ν_{2} and ν_{5} bands and l-type interactions in the ν_{5} band. The transition dipole moments squared were determined for each line, allowing determination of the vibrational transition moment and the Herman-Wallis coefficients of each band using the Aliev^{1} and Watson^{2} models. Comparison with data available in the literature permits testing the validity of these models applied to the ν_{2}/ν_{5} dyad system, strongly affected by resonances.

1MR. Aliev, D. Papousek and S. Urban, J. Mol. Spectrosc, 124, 285-325, (1987)

2JKG. Watson, J. Mol. Spectrosc, 153, 211-224, (1992)

p-number: p019  Submitted on Tue, 30 May 2023 14:02:26 +0200